OpenStudy (anonymous):

Ill give medals to whoever helps me... Solve the trinomal eqaution r^2-18r+56=0 ...clueless.........

OpenStudy (anonymous):

x=4 and x=14

OpenStudy (anonymous):

Why.?

OpenStudy (anonymous):

Factor the left hand side to (r-4)(r-14)=0 and use the zero product property to solve. Hence solve r-4=0 and r-14=0.